OpenStudy (anonymous):

At what speed v will the mass of an object be 10% greater than its rest mass?

OpenStudy (anonymous):

\[ \small{ \text{"It is not good to introduce the concept of the mass} \ M=m/ \sqrt{ 1- v^2/c^2} \\ \text{of a moving body for which no clear definition can be given. It is better to} \\ \text{introduce no other mass concept than the 'rest mass' } m \text{. Instead of introduc-} \\ \text{ing } M \text{ , it is better to mention the expression for the momentum and energy} \\ \text{ of a body in motion." } \\ \hspace{50px} \text{-Einstein } } \] Given the total energy of a an object in motion \[E=\gamma mc^2 \quad \quad ; \quad \quad \gamma=\frac{1}{\sqrt{1-\frac{v^2}{c^2}}}\] where its rest energy is at v=0 \[E_0 = mc^2\] we can see that total energy goes directly with mass; therefore, the question can be reframed as "At what speed v will the total energy of the object be 10% greater than the rest energy?" \[E=1.1E_0\] \[\gamma mc^2 = 1.1 (mc^2)\] \[\frac{1}{\sqrt{1-\frac{v^2}{c^2}}}=1.1\] Solve for v

OpenStudy (vincent-lyon.fr):

In France, we stick to Einstein's view, we have no other mass concept than 'rest mass' in our textbooks.